A chemical reaction takes place inside a flask submerged in a water bath. The water bath contains 3.30 kg of water at 22.9 degrees celcius. During the reaction 112. kJ of heat flows out of the bath and into the flask.

Calculate the new temperature of the water bath. You can assume the specific heat capacity of water under these conditions is 4.18·J·g−1K−1. Round your answer to 3 significant figures
Answer #1

When heat energy is added or released by the substance, the temperature will accordingly change. The relation between heat energy and temperature is different for different materials and it is specific heat that describes how they are related .

Now,

Q=m*c*∆t

Q = 112 * 10 ^3 joule

Mass = 3.30 kg = 3.30 *10^3 grams

C = 4.18

Putting in formula we get

∆t = q/m*c

∆t = 112 * 10 ^3 /3.30*10^3*4.18

∆t = 8.1194

∆ t = T1-T2 =8.1194

T2 = 22.9 - 8.1194

T2 = 14.7806°c